OpenStudy (anonymous):

Describe how the graph of y= x^2 can be transformed to the graph of the given equation. y = (x + 9)^2 – 3

OpenStudy (anonymous):

Shift the graph of y = x^2 left 9 units and then down 3 units. Shift the graph of y = x^2 down 9 units and then left 3 units. Shift the graph of y = x^2 right 9 units and then down 3 units. Shift the graph of y = x^2 up 9 units and then right 3 units.

OpenStudy (anonymous):

when i try to see the graph for y = (x + 9)^2 – 3 nothing comes up?

OpenStudy (anonymous):

does this helps

OpenStudy (anonymous):

?you guys send me 2 different graphs which one is right

OpenStudy (anonymous):

both are correct!

OpenStudy (anonymous):

just scale is different

OpenStudy (anonymous):

is it A?

OpenStudy (anonymous):

yes correct A

OpenStudy (anonymous):

but at @nickhouraney how did you get that graph

OpenStudy (anonymous):

i put in google y = (x + 9)^2 – 3 and got nothing

OpenStudy (anonymous):

@sara1234 just put this y=(x+9)^2-3 in google and press Enter key

OpenStudy (anonymous):

oh okay i understand lol just because i put spaces thanks @sami-21

OpenStudy (anonymous):

you 're welcome :)